OpenStudy (anonymous):

preform the indicated operation and simplify: x^2+3x-4 / 5x-5 TIMES 10x^2-40x / x^2 -16

OpenStudy (anonymous):

Work shown please for metal!!!!

OpenStudy (zehanz):

It looks like this:\[\frac{ x^2+3x-4 }{ 5x-5 }\cdot \frac{ 10x^2-40x }{ x^2-16 }\]Before doing anything else, it is always a good idea to completely factor everything:\[\frac{ (x-1)(x+4) }{ 5(x-1) }\cdot \frac{ 10(x+2)(x-2) }{ (x+4)(x-4) }\]You see the advantage of this: the fractions can be simplified before the multiplication is carried out:\[\frac{ x+4 }{ 5 }\cdot \frac{ 10(x+2)(x-2) }{ (x+4)(x-4) }=\frac{ 2(x+2)(x-2) }{ x-4 }\]You could expand now, but mostly is is better to leave it this way, because the zeros of numerator and denominator are in full sight. Conclusion: simplification can often be done before multiplication!